OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2, solve for g(f(3)). 5 7 9 10

OpenStudy (anonymous):

@Mehek14 would it be c?

OpenStudy (anonymous):

Here's what you do

OpenStudy (anonymous):

Substitute 3 for x in f(x) first

Mehek (mehek14):

\(f(x)=2*3+1\)

Mehek (mehek14):

multiply 2*3 first

OpenStudy (anonymous):

6+1 = 7

Mehek (mehek14):

correct so now put 7 for x in g(x)

Mehek (mehek14):

\(\dfrac{3*7-1}{2}\)

OpenStudy (anonymous):

so 10 would be the answer ;)

Mehek (mehek14):

yes :D

OpenStudy (anonymous):

you the besttt

Mehek (mehek14):

tyyyy :D